OpenStudy (anonymous):

to convert 23,000 seconds into hours which two ratios could you multiply by ?

OpenStudy (mathmale):

In other words, which two conversion factors could you multiply 23000 seconds by? How would you convert from seconds to minutes? from minutest o hours?

OpenStudy (anonymous):

seconds to hours

OpenStudy (mrnood):

you may find it easier to follow @mathmale it asks you for TWO ratios. Convert from seconds to minutes first, THEN from minutes to hours... \[seconds \div \frac{ seconds }{ \min } \div \frac{ \min }{ hour } = hours\]

OpenStudy (anonymous):

devided by 3600 23000/3600=230/36=6.38889 hrs.

OpenStudy (anonymous):

multiplied by 1/3600

OpenStudy (mathmale):

Note: 60 seconds = 1 minute, so the seconds-to-minutes conversion factor is \[\frac{ 1~\min }{ 60~\sec }\] I'd like to suggest that you multiply 23000 sec by this conversion factor. Your result will be the equivalent time in minutes instead of in seconds. Note: 60 minutes = 1 hour, so the minutes to hours conversion factor is ... what? What's the next step in the process of converting 23000 sec into hours?

OpenStudy (mathmale):

Note: It's also true that 3600 sec = 1 hour, so you could create just one conversion factor and then by using it, go immediately from 23000 sec to the equivalent number of hours.

OpenStudy (anonymous):

hey hey hey say directly my answer is best

OpenStudy (mrnood):

@kritikachandak It is totally evident that your answer is the most direct way to convert seconds to hours. However, it is NOT the best answer BECAUSE the question asks explicitly for TWO ratios. Therefore your solution does not answer the question at all.